Circular No. 9129

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET P/2001 Q11 (NEAT) Maik Meyer, Limburg, Germany, reports his discovery of a comet on

images obtained by the NEAT project on three nights at Palomar and on five nights at Haleakala during Aug.–Dec. 2001. The 2001 Aug. 18 Palomar discovery-night images (discovery observation tabulated below) show a 21′′ tail in p.a. 256^{o}, and there was a 0′.2 tail in p.a. 264^{o} on the Aug. 22 Haleakala frames. The Aug. 18 observations were reported (though not as being cometary) by NEAT at the time (mag 18.6–18.7), but the object was never followed up; likewise for four LONEOS observations (mag 17.9) on Oct. 24.

2001 UT α_{2000} δ_{2000} Mag. Observer
Aug. 18.47017 3^{h}18^{m}48.^{s}31 −2^{o}37′06′′.8 18.9 NEAT

The detailed astrometry, an elliptical orbit (epoch 2001 June 20.0 TT, T = 2001 June 22.3510 TT, q = 1.851629 AU, e = 0.450257, ω = 202^{o}.2729, Ω = 147^{o}.0124, i = 20^{o}.8526, equinox 2000.0, P = 6.18 years), and residuals are given on MPEC 2010-F38. Searches of relevant frames by Meyer and others have so far failed to show the comet at its return in 2007.

PRINTED BACK ISSUES Since its move to Cambridge in 1965, the Central Bureau has main-

tained an extensive archive of back issues of the printed IAUCs; these have been sent to interested individuals upon request (for a small charge). Having inherited some years’ worth of back issues from Copenhagen Observatory at the time of the 1965 transfer, we now have several fairly continuous sets of printed IAUCs going back to No. 1727 (dated 1960 May 21). Unfortunately, the Bureau cannot continue to maintain such a large collection of back issues, which now take up several large card-file cabinets. Therefore, before discarding any of the back issues of these Circulars, we now ask that readers and libraries who might be interested in obtaining sets of the cards — to extend their existing collections — should contact the undersigned as soon as possible to make such a request.

CORRIGENDA On IAUC 9123, ‘COMET C/2010 DG_{56} (WISE)’, the second MPEC

reference should read 2010-E42, not 2010-E43. On IAUC 9127, line 13, for Campinas, Brazil, read Porto Seguro, Brazil,

2010 March 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green